# Diagonalization of Polynomial-Time Deterministic Turing Machines via Nondeterministic Turing Machines

### v33

@@ Abstract.
The diagonalization technique was invented by Georg Cantor to show that there are more real numbers than algebraic numbers and is very important in theoretical computer science. In this work, we enumerate all of the polynomial-time deterministic Turing machines and diagonalize against all of them by a universal nondeterministic Turing machine. As a result, we obtain that there is a language $L_{d}$ not accepted by any polynomial-time deterministic Turing machines but accepted by a nondeterministic Turing machine running within time $O(n^{k})$ for any $k\in\mathbb{N}_{1}$. Based on these, we further show that $L_{d}\in\mathcal{NP}$. That is, we present a proof that $\mathcal{P}$ and $\mathcal{NP}$ differ. Meanwhile, we show that there exists a language $L_{s}$ in $\mathcal{P}$, but the machine accepting it also runs within time $O(n^{k})$ for all $k\in\mathbb{N}_{1}$. Lastly, we show that if $\mathcal{P}^{O}=\mathcal{NP}^{O}$ and on some rational base assumptions, then the set $P^{O}$ of all polynomial-time deterministic oracle Turing machines with oracle $O$ is not enumerable, thus demonstrating that the diagonalization technique (via a universal nondeterministic oracle Turing machine) will generally not apply to the relativized versions of the $\mathcal{P}$ versus $\mathcal{NP}$ problem.
@@ 1. Introduction
More interestingly, the technique of diagonalization led to some early successes in computational complexity theory. In history, it was invented by Georg Cantor [Can91] to show that there are more real numbers than algebraic numbers, then it was used by Alan Turing to show that there exists a sequence that is not computable (see [Tur37], p. 246), and then refined to prove computational complexity lower bounds by Hartmanis and Stearns [HS65]. A typical theorem in this area is that more time/space buys more computational power [HS65, Coo73, FS07, Pap94, Sip13, AB09]. For instance, there are functions computable in time $n^{2}$, say, which are not computable in time $n$. At present, we assume that the heart of such arguments is the existence of a universal Turing machine, which can simulate every other Turing machine with only a small loss in efficiency.111 At the end of the work, it will become clear that the notion of Enumerable of Turing machines is an important prerequisite for application of diagonalization.
As we mentioned earlier, the diagonalization technique was invented by Cantor [Can91] to show that there are more real numbers than algebraic numbers, and early it demonstrated powers in separating two complexity classes. There is an important doubt: can the diagonalization approach resolve the $\mathcal{P}$ versus $\mathcal{NP}$ problem? Researchers realized in the $1970$s that diagonalization alone may not be able to resolve the $\mathcal{P}$ versus $\mathcal{NP}$ problem; it needs other new techniques besides diagonalization (see e.g., [Sip13, AB09]). But, at the end of the work, we will say that the above point of view is partially correct and leave aside why we think so.
@@ 1.1. Main Results
We put our aforementioned curiosities, or ideas, into practice in this work. We first enumerate all polynomial-time deterministic Turing machines222 It is much more suitable to say that we first present a way to enumerate all of the polynomial-time deterministic Turing machines. and then diagonalize against all of them by a universal nondeterministic Turing machine. Generally, complexity theorists used universal deterministic Turing machines to diagonalize against a list of deterministic Turing machines (see e.g., [HS65, Pap94, AB09]), or applied universal nondeterministic Turing machines to diagonalize against a list of nondeterministic Turing machines; see e.g., [AB09, For00, FS07, Zak83]. Using a universal nondeterministic Turing machine to diagonalize against a collection of deterministic Turing machines appears less often in the previous literature, except in the author’s recent work [Lin21], and it is maybe a new attempt that may lead to success. As an amazing result, we obtain the following important theorem:
@@
In addition to the “Relativization Barrier,” some theorists in relativized worlds also developed the so-called “Algebrization Barrier”; see [AW09]. Also, the reader can find some interesting discussions about these barriers in the reference [AB18].
Now let us return to the proof technique of diagonalization once again. Cantor’s diagonal process, also called the diagonalization argument, was published in $1891$ by Georg Cantor [Can91] as a mathematical proof that there are infinite sets that cannot be put into one-to-one correspondence with the infinite set of positive numbers, i.e., $\mathbb{N}_{1}$ defined in the following Section 2. The technique of diagonalization was first used in computability theory in the $1930$s by Turing [Tur37] to show that there exists undecidable language. In computational complexity theory, in their seminal paper [HS65], Hartmanis and Stearns employed the diagonalization proof to give time hierarchy. For more other summarization about it, please consult [For00] for a survey. On the other hand, some complexity theorists or textbooks (see e.g., page $73$ in [AB09]) in computational complexity theory regard that “diagonalization” is any technique that relies solely upon the following properties of Turing machines:
I:
The existence of an effective representation of Turing machines by strings.
II:
The ability of one Turing machine simulate any other without much overhead in running time or space.
Thus, the computational complexity textbook [AB09] thinks of these properties as also applicable for oracle Turing machines and further regards that to show $\mathcal{P}\neq\mathcal{NP}$ requires some other properties in addition to the properties I and II as stated above. However, we would like to say that this kind of perspective is not absolutely (fully) correct either. Notice also that a similar point of view was followed in [Sip13, BC94], i.e., the references [Sip13] and [BC94] hold the viewpoint that if we could prove that $\mathcal{P}$ and $\mathcal{NP}$ were different by diagonalizing, then we could conclude that they are different relative to any oracle as well. Interestingly enough, we will demonstrate the key points besides the aforementioned assumptions I and II when diagonalization techniques are applicable such that we can overcome the so-called “Relativization Barrier.”
But at present, for the convenience of the reader, let us first quote the notion of an enumeration of a set as follows:
@@ 1.2. Our Approach
As a novel idea and attempt, we enumerate all polynomial-time deterministic Turing machines and then diagonalize against all of them with a universal nondeterministic Turing machine, thus obtaining a language $L_{d}$ not in $\mathcal{P}$. Based on the above work, we then carefully do an analysis for the language $L_{d}$, showing that this language is in $\mathcal{NP}$. To stress that it is “novel” is because, in general, a common practice is to use a universal deterministic Turing machine to diagonalize against a list of deterministic Turing machines (see e.g., [Pap94, AB09]), or to employ a universal nondeterministic Turing machine to diagonalize against a list of nondeterministic Turing machines; see e.g., [AB09, For00, FS07, Zak83]. That using a universal nondeterministic Turing machine to diagonalize against all of the polynomial-time deterministic Turing machines is less common and can be seen as a new attempt. We would like to emphasize that this approach has seldom been tried before in the literature, except that recently we noted that in [Sip92], Sipser mentioned an idea (but this simple argument fails; see [Sip92], p. 605) to give a nondeterministic polynomial-time Turing machine that has an opportunity to run each of the deterministic polynomial-time Turing machines and arrange to accept a differing language.
@@ 2. Preliminaries
Let $\mathbb{N}$ denote the natural numbers
@@ 4. Diagonalization against All Polynomial-Time Deterministic Turing Machines
The nondeterministic Turing machine $M_{0}$ described above is of time complexity, say $S$, which is currently unknown. According to Lemma 10, $M_{0}$ is equivalent to a single-tape nondeterministic $O(S^{2})$ time-bounded Turing machine, and it of course accepts some language $L_{d}$.
@@
holds or not.
@@
then to simulate $N$ and accept the input if and only if $N$ accepts; otherwise, $M_{0}$ rejects the input. Note that such a design does not change $M_{0}$’s time complexity.
@@
as observed by Cook [Coo73]. Thus, in his work [Coo73], Cook uses different techniques and then, by “indirect” diagonalization, which is somewhat complicated (e.g., the ingredient of technique from [Iba72] with other complicated techniques), to show a nondeterministic time hierarchy theorem [Coo73] (it has to say that there exist other techniques such as lazy diagonalization that are also capable of showing nondeterministic time hierarchy theorems; see e.g., [AB09, For00, FS07, Zak83]). But fortunately, we diagonalize against all polynomial-time deterministic Turing machines rather than polynomial-time nondeterministic Turing machines in Theorem 4.1 and its proof, so our nondeterministic Turing machine $M_{0}$ can flip the answer immediately when the simulation ends.
@@ 5.1. The $2$nd Simpler Proof
Originally, we call $M_{0}$ a polynomial-time nondeterministic Turing machine. But complexity theorist Lance Fortnow [For21]252525 We received valuable criticisms [For21] from Lance Fortnow after posting the manuscript of this work on the arXiv. argues that $M_{0}$ does not run in polynomial time because it runs in time $O(n^{k})$ for any $k\in\mathbb{N}_{1}$. He thinks the notion of a polynomial-time machine is a fixed mathematical definition, i.e., call a machine running in polynomial time iff the machine runs in time $O(n^{c})$ for some fixed constant $c>0$. For this, see special thanks expressed in acknowledgements subsection LABEL:sec:acks.
@@ 6.2. Proof of $L_{s}$ in $\mathcal{P}$
Can we now deny that the language $L_{s}$ is not in $\mathcal{P}$ due to the machine $M^{\prime}_{0}$ accepting it running within $O(n^{k})$ for all $k\in\mathbb{N}_{1}$? Of course, we can not, because the language $L_{s}$ accepted by $M^{\prime}_{0}$ is in $\mathcal{P}$ but not others.282828 In fact, we can construct a universal deterministic Turing machine running within time $O(n^{i})$ for any $i\in\mathbb{N}_{1}$ accepting the language $L^{\prime}_{s}=\{(\langle M_{i}\rangle,w)\,|\,w\in L(M_{i})\},$ where $\langle M_{i}\rangle$ is the binary string representation of some polynomial-time deterministic Turing machine taken from the set $\{(M,k)\}$, and $L(M_{i})$ denotes the language accepted by $M_{i}$, and the tuple of $(\langle M_{i}\rangle,w)$ can be seen as a pairing function, which is similar to [DK14]; see Proposition 1.17, p. 26 in [DK14]. Moreover, still denote the set of all the polynomial-time deterministic Turing machines by $\{(M,k)\}$, then
@@ 7.1. Proof of Theorem 6
time-bounded oracle Turing machine. Let $D^{O}_{i}$ have $s$ states and $t$ tape symbols. Since303030 See (III) of footnote $6$.
@@ 8. Concluding Remarks and Open Problems
Are these two complexity classes the same? Also note that there is a subfield of computational complexity theory, namely, proportional proof complexity, which was initiated by Cook and Reckhow [CR79] and is devoted to the goal of proving the conjecture

### v34

@@ Abstract.
The diagonalization technique was invented by Georg Cantor to show that there are more real numbers than algebraic numbers and is very crucial in theoretical computer science. In this work, we enumerate all of the polynomial-time deterministic Turing machines and diagonalize against all of them by a universal nondeterministic Turing machine. As a result, we obtain that there is a language $L_{d}$ not accepted by any polynomial-time deterministic Turing machines but accepted by a nondeterministic Turing machine running within time $O(n^{k})$ for any $k\in\mathbb{N}_{1}$. Based on these, we further show that $L_{d}\in\mathcal{NP}$. That is, in this work, we present a proof that $\mathcal{P}$ and $\mathcal{NP}$ differ. Meanwhile, we show that there exists a language $L_{s}$ in $\mathcal{P}$, but the machine accepting it also runs within time $O(n^{k})$ for all $k\in\mathbb{N}_{1}$. Lastly, we show that if $\mathcal{P}^{O}=\mathcal{NP}^{O}$ and on some rational base assumptions, then the set $P^{O}$ of all polynomial-time deterministic oracle Turing machines with oracle $O$ is not enumerable, thus demonstrating that the diagonalization technique (via a universal nondeterministic oracle Turing machine) will generally not apply to the relativized versions of the $\mathcal{P}$ versus $\mathcal{NP}$ problem.
@@ 1. Introduction
More interestingly, the technique of diagonalization led to some early successes in computational complexity theory. In history, it was invented by Georg Cantor [Can91] to show that there are more real numbers than algebraic numbers, then it was used by Alan Turing to show that there exists a sequence that is not computable (see [Tur37], p. 246), and then refined to prove computational complexity lower bounds by Hartmanis and Stearns [HS65]. A typical theorem in this area is that more time/space buys more computational power [HS65, Coo73, FS07, FS17, Pap94, Sip13, AB09]. For instance, there are functions computable in time $n^{2}$, say, which are not computable in time $n$. At present, we assume that the heart of such arguments is the existence of a universal Turing machine, which can simulate every other Turing machine with only a small loss in efficiency. 111 At the end of the work, it will become clear that the notion of Enumerable of Turing machines is an important prerequisite for application of diagonalization techniques.
As we mentioned earlier, the diagonalization technique was invented by Cantor [Can91] to show that there are more real numbers than algebraic numbers, and early on it demonstrated powers in separating two complexity classes. There is an important doubt: can the diagonalization approach resolve the $\mathcal{P}$ versus $\mathcal{NP}$ problem? Researchers realized in the $1970$s that diagonalization alone may not be able to resolve the $\mathcal{P}$ versus $\mathcal{NP}$ problem; it needs other new techniques besides diagonalization (see e.g., [Sip13, AB09]). But, at the end of the work, we will say that the above point of view is partially correct and leave aside why we think so.
@@ 1.1. Main Results
We put our aforementioned curiosities, or ideas, into practice in this work. We first enumerate all polynomial-time deterministic Turing machines and then diagonalize against all of them by a universal nondeterministic Turing machine.222 It is much more suitable to say that we first present a way to enumerate all of the polynomial-time deterministic Turing machines. Generally, complexity theorists used universal deterministic Turing machines to diagonalize against a list of deterministic Turing machines (see e.g., [HS65, Pap94, AB09]), or applied universal nondeterministic Turing machines to diagonalize against a list of nondeterministic Turing machines; see e.g., [AB09, For00, FS07, FS17, Zak83]. Using a universal nondeterministic Turing machine to diagonalize against a collection of deterministic Turing machines appears less often in the previous literature, except in the author’s recent work [Lin21], and it is maybe a new attempt that may lead to success. As an amazing result, we obtain the following important theorem:
@@
In addition to the “Relativization Barrier,” some theorists in relativized worlds also developed the so-called “Algebrization Barrier;” See [AW09]. Also, the reader can find some interesting discussions about these barriers in the reference [AB18].
Now let us return to the proof techniques of diagonalization once again. Cantor’s diagonal process, also called the diagonalization argument, was published in $1891$ by Georg Cantor [Can91] as a mathematical proof that there are infinite sets that cannot be put into one-to-one correspondence with the infinite set of positive numbers, i.e., $\mathbb{N}_{1}$ defined in the following Section 2. The technique of diagonalization was first used in computability theory in the $1930$s by Turing [Tur37] to show that there exists undecidable language. In computational complexity theory, in their seminal paper [HS65], Hartmanis and Stearns employed the diagonalization proof to give time hierarchy. For more other summarization about it, please consult [For00] for a survey. On the other hand, Arora and Barak [AB09] (see e.g., page $73$ in [AB09], which is a textbook in computational complexity theory) regard that “diagonalization” is any technique that relies solely upon the following properties of Turing machines:
I: The existence of an effective representation of Turing machines by strings.
II: The ability of one Turing machine simulate any other without much overhead in running time or space.
Thus, Arora and Barak, the authors of the computational complexity textbook [AB09], and other complexity theorists think of these properties as also applicable for oracle Turing machines and further regard that to show $\mathcal{P}\neq\mathcal{NP}$ requires some other properties in addition to the properties I and II as stated above. However, we would like to say that this kind of perspective is not absolutely (fully) correct either. Notice also that a similar point of view was followed in [Sip13, BC94], i.e., the references [Sip13] and [BC94] hold the same viewpoint that if we could prove that $\mathcal{P}$ and $\mathcal{NP}$ were different by diagonalizing, then we could conclude that they are different relative to any oracle as well. Interestingly enough, we will demonstrate the key points besides the aforementioned assumptions I and II when diagonalization techniques are applicable such that we can overcome the so-called “Relativization Barrier.”
But at the moment, for the convenience of the reader, let us first quote a mathematical definition of an enumeration of a set, which appears in modern mathematics textbooks such as [Rud76], as follows:
@@ 1.2. Our Approach
As a novel idea and attempt, we enumerate all polynomial-time deterministic Turing machines and then diagonalize against all of them with a universal nondeterministic Turing machine, thus obtaining a language $L_{d}$ not in $\mathcal{P}$. Based on the above work, we then carefully do an analysis for the language $L_{d}$, showing that this language is in $\mathcal{NP}$. To stress that it is “novel” is because, in general, a common practice is to use a universal deterministic Turing machine to diagonalize against a list of deterministic Turing machines (see e.g., [Pap94, AB09]), or to employ a universal nondeterministic Turing machine to diagonalize against a list of nondeterministic Turing machines; see e.g., [AB09, For00, FS07, FS17, Zak83]. That using a universal nondeterministic Turing machine to diagonalize against all of the polynomial-time deterministic Turing machines is less common and can be seen as a new attempt. We would like to emphasize that this approach has seldom been tried before in the literature, except that recently we noted that in [Sip92], Sipser mentioned an idea (but this simple argument fails; see [Sip92], p. 605) to give a nondeterministic polynomial-time Turing machine that has an opportunity to run each of the deterministic polynomial-time Turing machines and arrange to accept a differing language.
@@ 2. Preliminaries
Let $\mathbb{N}$ denote the set of natural numbers
@@ 4. Diagonalization against All Polynomial-Time Deterministic Turing Machines
The nondeterministic Turing machine $M_{0}$ described above is of time complexity, say $S$, which is currently unknown. According to Lemma 10, $M_{0}$ is equivalent to a single-tape nondeterministic
$O(S^{2})$
time-bounded Turing machine, and it of course accepts some language $L_{d}$.
@@
holds or not. Currently, we completely have no idea about how to prove or disprove this important relationship.
@@
then simulate $N$ nondeterministically in time $n^{k+1}$ and output its answer, i.e., accepting if $N$ accepts and rejecting if $N$ rejects; otherwise, $M_{0}$ rejects the input. Note that such a design does not change $M_{0}$’s time complexity.
@@
as observed by Cook [Coo73]. Thus, in his work [Coo73], Cook uses different techniques and then, by “indirect” diagonalization, which is somewhat complicated (e.g., the ingredient of technique from [Iba72] with other complicated techniques), to show a nondeterministic time hierarchy theorem [Coo73] (it has to say that there exist other techniques such as lazy diagonalization that are also capable of showing nondeterministic time hierarchy theorems; see e.g., [AB09, For00, FS07, Zak83]. More surprisingly, Fortnow [FS17] developed a much more elegant and simple style of diagonalization to show nondeterministic time hierarchy). But fortunately, we diagonalize against all polynomial-time deterministic Turing machines rather than polynomial-time nondeterministic Turing machines in Theorem 4.1 and its proof, so our nondeterministic Turing machine $M_{0}$ can flip the answer immediately when the simulation ends.
@@ 5.1. The $2$nd Simpler Proof
Originally, we call $M_{0}$ a polynomial-time nondeterministic Turing machine. But complexity theorist Lance Fortnow [For21]252525 We received valuable criticisms [For21] from Lance Fortnow shortly after posting the manuscript of this work on the arXiv. argues that $M_{0}$ does not run in polynomial time because it runs in time $O(n^{k})$ for any $k\in\mathbb{N}_{1}$. He thinks the notion of a polynomial-time machine is a fixed mathematical definition, i.e., call a machine running in polynomial time iff the machine runs in time $O(n^{c})$ for some fixed constant $c>0$. For this, see special thanks expressed in the acknowledgements section; see Section LABEL:sec:acks.
@@ 6.2. Proof of $L_{s}$ in $\mathcal{P}$
Can we now deny that the language $L_{s}$ is not in $\mathcal{P}$ due to the machine $M^{\prime}_{0}$ accepting it running within time
$O(n^{k})$
for all $k\in\mathbb{N}_{1}$? Of course, we can not, because the language $L_{s}$ accepted by $M^{\prime}_{0}$ is in $\mathcal{P}$ but not others.282828 In fact, we can construct a universal deterministic Turing machine running within time $O(n^{i})$ for any $i\in\mathbb{N}_{1}$ accepting the language $L^{\prime}_{s}=\{(\langle M_{i}\rangle,w)\,|\,w\in L(M_{i})\},$ where $\langle M_{i}\rangle$ is the binary string representation of some polynomial-time deterministic Turing machine taken from the set $\{(M,k)\}$, and $L(M_{i})$ denotes the language accepted by $M_{i}$, and the tuple of $(\langle M_{i}\rangle,w)$ can be seen as a pairing function, which is similar to [DK14]; see Proposition 1.17, p. 26 in [DK14]. Moreover, still denote the set of all the polynomial-time deterministic Turing machines by $\{(M,k)\}$, then
@@ 7.1. Proof of Theorem 6
time-bounded oracle Turing machine. Let $D^{O}_{i}$ have $s$ states and $t$ tape symbols. Since $M^{O}_{0}$’s simulation can be done within time
$O(T(n)\log T(n)),$
we thus have that
@@ 8. Concluding Remarks and Open Problems
Namely, are these two complexity classes the same? Also note that there is a subfield of computational complexity theory, namely, proportional proof complexity, which was initiated by Cook and Reckhow [CR79] and is devoted to the goal of proving the conjecture